OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 3?

OpenStudy (anonymous):

an = 5 − 3(n − 1); all integers where n ≥ 1 an = 5 − 3(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 1